±√324 - how do you solve this
Answer by tommyt3rd(5050)

A method of extracting roots is to use the factorization of a number (by successively dividing out primes until the number is manageable)
factoring 324...
324=2*162=
2^2*81=2^2*9^2

so that 324 = (2*9)^2 = 18^2

therefore the square root of 324 is 18
±√324 = ±18